# Probing Mechanical Reasoning in Large Vision Language Models

### v1

@@ 3.3 VLMs Understand Gear Systems
3.4 Seesaw-like systems and Leverage Principle
@@
We observe diverse behaviors in VLMs on solving seesaw-like mechanical systems and applying leverage principle (Figure 5). Experiments 5A, 5C, 5D, and 5F involve relatively simple scene-based question answering and basic extrapolation. Experiments 5A and 5D are straightforward question-answer tasks. As long as the model identifies the key information in the images—such as the length of the wrench in Experiment 5A or the unequal torques in Experiment 5D—it can easily solve the problems. Experiments 5C and 5F involve simple extrapolations based on scene information, like understanding how changes in the length of the resistance arm affect force in Experiment 5C or predicting the outcome when the weights on a balance scale are unequal in Experiment 5F. It is evident that the model performed exceptionally well in these experiments, providing both correct answers and detailed explanations. However, in Experiments 5B and 5E, VLMs failed to give the correct answers. We suspect that VLMs struggle when the mechanical reasoning processes, particularly applying the leverage principle and solving seesaw-like systems, require multiple steps. For example, in Experiment 5B, while VLMs recognize it is a seesaw-like system, they fail to understand how to operate within this system: sliding backward would give more leverage. This reasoning process may require backtracking through several steps: first, identifying the seesaw-like system; second, understanding how to apply leverage; third, realizing that sliding backward increases leverage; fourth, recognizing that more leverage is needed to level the system; and finally, concluding that sliding backward is necessary. The same issue arises in Experiment 5E.
3.5 Inertia and Motion
@@
Experiment 6E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 6A, 6B, and 6D, on the other hand, are cases of simple extrapolation. They are categorized as "simple extrapolation" because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 6A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 6B and 6D.
Experiments 6C and 6F, however, involve more complex reasoning related to inertia and motion prediction. Experiment 6F is a derivative question from Experiment 6E, asking what would happen if the cart suddenly stopped. The suspended object would continue moving forward due to inertia. Although the scene in Experiment 6F is similar to Experiment 6E, and the question is similar to that in Experiment 6A, it involves more objects (the pulling rope, the cart, and the pulled object), more physical principles, and a sudden change in force (as the rope loses tension). The model clearly struggled with this problem, providing an incorrect answer and an explanation that did not meet expectations. Notably, Experiment 6C asks a common-sense question based on pendulum motion. The image in Experiment 6C marks three points in the half-arc of the pendulum’s swing: the highest point, the quarter-arc point, and the lowest point. This information is also specified in the prompt. However, the model gave an incorrect answer, and its explanation was confused. The model completely misunderstood the designated points in the diagram and in the prompt, mistaking the first and third points as the left and right highest points of the pendulum’s motion and the second point as the lowest point. One possible reason for this error is that the model may have mistaken the half-arc pendulum motion in Experiment 6C for a full pendulum motion. A deeper explanation might suggest that the model has issues with perceptual constancy. The image in Experiment 6C might have been misinterpreted as a 3D scene, and this visual misperception could be the root of the error. This hypothesis requires further experimentation to be confirmed.
3.6 Fluid Systems
@@
Experiments 7B and 7C highlighted the model’s weaker inference abilities. In Experiment 7B, the hole in the bucket was positioned at the middle of the bucket’s wall. Clearly, once the water level drops to the level of the hole, water will stop flowing out, meaning some water will remain in the bucket. However, the model failed to capture this crucial detail about the hole’s location, leading to an incorrect answer. Similarly, in Experiment 7C, although the model provided the correct answer, its explanation was incorrect.
@@
It is worth noting that the model excelled in image comprehension and detail recognition, especially in Experiments 7D and 7F. In Experiment 7D, the measuring cup had a narrow base and a wider top, meaning the scale markings could not be evenly spaced. The model successfully captured these details, including the design of the measuring cup, the distribution of the scale, and the corresponding numerical values and units, offering a detailed explanation. The same can be said for Experiment 7F. The model accurately identified that the balloon was fully inflated with gas and incorporated the concepts of gas and liquid density to explain the current state and predict the next step of the experiment.
4 Discussion
Here, we leverage MechBench, an assay of 141 cognitive experiments that include many aspects of mechanical reasoning to investigate Vision Language Models. We investigate VLM abilities in reasoning about mechanical stability, gear systems, pulley systems, inertia and motion, and fluid systems. In general, we observe very diverse phenomena in VLM mechanical reasoning abilities. It’s definitely the case that VLMs already have some sorts of mechanical reasoning abilities. Nevertheless, we also observe that VLMs have numerous failure modes in reasoning about physical objects and their mechanical interactions in the world.
One parallelism might be drawn here between current machine intelligence and animal intelligence. Whether animals have the ability to mechanically reason about objects in the world still remain a hotly debated scientific topic [Hall, 1963, Van Lawick-Goodall, 1971, Seed and Byrne, 2010, Andrews, 2020]. Some examples of mechanical reasoning in animals have been observed. Chimpanzees, for example, could transform a leafy vine into a thin, flexible tool for fishing termites, while crafting a sturdier rod for dipping into ant nests [Goodall, 1986]. A New Caledonian crow uses its beak to shape the serrated leaves of a Pandanus tree into strips, which the crow then employs to extract insects or larvae from hard-to-reach places [St Clair and Rutz, 2013]. Nevertheless, animals are not able to achieve the level of perplexity and creativity in terms of manipulating and building mechanical systems comparing to humans [Birch et al., 2020]. One interesting direction would be juxtaposing animal intelligence, machine intelligence, and human intelligence with our mechanical reasoning assay, and learn from the similarities and differences in how animals, machines, and humans reason about physical objects and their interactions in the world.
The computational basis of mechanical reasoning in the human mind has been a focus of research. One of the core theories in cognitive science regarding human mechanical reasoning is mental simulation. As proposed by Hegarty [1992], when people reason about mechanical systems, such as gears or pulley systems, they engage in a step-by-step mental process where each component is animated sequentially. This is different from imagining a static scene; rather, people simulate the dynamic changes within the system incrementally. This mechanism reflects a fundamental limitation in working memory, in that humans cannot animate the entire system holistically, but instead rely on breaking it down into smaller causal units to reason through the interaction of its components.
Similarly, in our experiments, Vision-Language Models (VLMs) exhibited signs of piecemeal reasoning but showed limitations in comprehending the full causal chain, much like the sequential cognitive processes seen in human mental simulation. The VLMs often struggled to reason about components deeper within the causal chain, such as inferring the behavior of a secondary gear or pulley based on the motion of a primary one. This suggests that while VLMs can engage with the initial stages of a problem, they may not yet possess the cognitive flexibility required to simulate complex mechanical systems across multiple stages.
Furthermore, studies by [Schwartz and Black, 1996] demonstrated that humans use analog processes in mechanical reasoning, meaning their cognitive processes mirror the physical dynamics they aim to understand. For example, when reasoning about the interaction between gears, humans often rely on mental rotation—where the time to mentally rotate an object is proportional to the angle of rotation. This ability to accurately simulate physical processes is crucial for reasoning about real-world mechanical systems, such as predicting how two interlocking gears will turn.
However, our analysis of VLMs in the gear systems tasks from MechBench reveals that while VLMs sometimes mimic human-like performance on surface-level tasks, their ability to engage in these analog processes remains underdeveloped. VLMs often fail to understand proportionality in systems, such as recognizing how changes in one gear affect subsequent ones. This suggests that their internal representation of mechanical systems lacks the depth of human analog reasoning.
Finally, the role of spatial representation in mechanical reasoning is crucial. Dual-task studies by [Hegarty and Sims, 1994] revealed that spatial working memory plays a significant role in solving mechanical reasoning tasks, as visuospatial information needs to be continuously processed and manipulated. VLMs, while equipped with powerful vision-language integration, appear to lack this level of spatial manipulation, often failing in tasks that require maintaining a coherent internal spatial representation of the objects involved. This limitation leads to errors when reasoning about inertia, fluid dynamics, or the spatial consequences of motion within a system.
Thus, by applying insights from human mechanical reasoning research to our findings with VLMs, it is evident that current VLMs do not yet fully replicate the sophisticated cognitive processes humans use. Mechanical reasoning requires the integration of mental simulation, analog processes, and spatial manipulation, all of which present ongoing challenges for machine intelligence. The path forward involves enhancing VLMs’ ability to engage in these types of reasoning, which could be guided by further exploration of how humans—and even animals—approach the same problems.
Acknowledgments and Disclosure of Funding
The project is not funded by any organizations.

### v2

@@ A.3 Gear Systems
A.4 Seesaw-like systems and Leverage Principle
@@
We observe diverse behaviors in VLMs on solving seesaw-like mechanical systems and applying leverage principle (Figure 10). Experiments 10A, 10C, 10D, and 10F involve relatively simple scene-based question answering and basic extrapolation. Experiments 10A and 10D are straightforward question-answer tasks. As long as the model identifies the key information in the images—such as the length of the wrench in Experiment 10A or the unequal torques in Experiment 10D—it can easily solve the problems. Experiments 10C and 10F involve simple extrapolations based on scene information, like understanding how changes in the length of the resistance arm affect force in Experiment 10C or predicting the outcome when the weights on a balance scale are unequal in Experiment 10F. It is evident that the model performed exceptionally well in these experiments, providing both correct answers and detailed explanations. However, in Experiments 10B and 10E, VLMs failed to give the correct answers. We suspect that VLMs struggle when the mechanical reasoning processes, particularly applying the leverage principle and solving seesaw-like systems, require multiple steps. For example, in Experiment 10B, while VLMs recognize it is a seesaw-like system, they fail to understand how to operate within this system: sliding backward would give more leverage. This reasoning process may require backtracking through several steps: first, identifying the seesaw-like system; second, understanding how to apply leverage; third, realizing that sliding backward increases leverage; fourth, recognizing that more leverage is needed to level the system; and finally, concluding that sliding backward is necessary. The same issue arises in Experiment 5E.
A.5 Inertia and Motion
@@
Experiment 11E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 11A, 11B, and 11D, on the other hand, are cases of simple extrapolation. They are categorized as "simple extrapolation" because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 11A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 11B and 11D.
Experiments 11C and 11F, however, involve more complex reasoning related to inertia and motion prediction. Experiment 11F is a derivative question from Experiment 11E, asking what would happen if the cart suddenly stopped. The suspended object would continue moving forward due to inertia. Although the scene in Experiment 11F is similar to Experiment 11E, and the question is similar to that in Experiment 11A, it involves more objects (the pulling rope, the cart, and the pulled object), more physical principles, and a sudden change in force (as the rope loses tension). The model clearly struggled with this problem, providing an incorrect answer and an explanation that did not meet expectations. Notably, Experiment 11C asks a common-sense question based on pendulum motion. The image in Experiment 11C marks three points in the half-arc of the pendulum’s swing: the highest point, the quarter-arc point, and the lowest point. This information is also specified in the prompt. However, the model gave an incorrect answer, and its explanation was confused. The model completely misunderstood the designated points in the diagram and in the prompt, mistaking the first and third points as the left and right highest points of the pendulum’s motion and the second point as the lowest point. One possible reason for this error is that the model may have mistaken the half-arc pendulum motion in Experiment 11C for a full pendulum motion. A deeper explanation might suggest that the model has issues with perceptual constancy. The image in Experiment 11C might have been misinterpreted as a 3D scene, and this visual misperception could be the root of the error. This hypothesis requires further experimentation to be confirmed.
A.6 Fluid Mechanics
@@
Experiments 12B and 12C highlighted the model’s weaker inference abilities. In Experiment 12B, the hole in the bucket was positioned at the middle of the bucket’s wall. Clearly, once the water level drops to the level of the hole, water will stop flowing out, meaning some water will remain in the bucket. However, the model failed to capture this crucial detail about the hole’s location, leading to an incorrect answer. Similarly, in Experiment 12C, although the model provided the correct answer, its explanation was incorrect.
@@
It is worth noting that the model excelled in image comprehension and detail recognition, especially in Experiments 12D and 12F. In Experiment 12D, the measuring cup had a narrow base and a wider top, meaning the scale markings could not be evenly spaced. The model successfully captured these details, including the design of the measuring cup, the distribution of the scale, and the corresponding numerical values and units, offering a detailed explanation. The same can be said for Experiment 12F. The model accurately identified that the balloon was fully inflated with gas and incorporated the concepts of gas and liquid density to explain the current state and predict the next step of the experiment.